OpenStudy (anonymous):

List all the solutions of the equation |8w+4| = 68 .

OpenStudy (dumbcow):

set up 2 cases 8w+4 = 68 and 8w+4 = -68

OpenStudy (anonymous):

but how do i find W?

OpenStudy (dumbcow):

get it by itself on left hand side by canceling everything else ex: 8w+4 = 68 to cancel out the addition of 4 i subtract 4 on both sides 8w +4 = 68 -4 -4 _____________ 8w +0 = 64 cancel out multiplication of 8 by dividing by 8 on both sides 8w/8 = 64/8 w = 8 make sense?

OpenStudy (anonymous):

kind ok...so there are 2 solutions.... 8, -8 right?

OpenStudy (anonymous):

kind of*

OpenStudy (dumbcow):

not quite, its not always just the negative of the answer for first one remember there were 2 case follow same steps to find w again for 8w +4 = -68

OpenStudy (anonymous):

so -9, and 8?

OpenStudy (dumbcow):

yes correct :)